Find the resultant of the following


vectors.


A = 30 m at 30° with the x-axis


B = 20 m at 120° with the x-axis


C = 10 m at 225° with the x-axis

1
Expert's answer
2021-12-24T11:31:59-0500

Find the components of the vectors:


"A_x = 30\cos30°,\\
B_x = 20\cos120°,\\
C_x = 10\cos225°,\\
R_x=A_x+B_x+C_x=8.9\text{ m}.\\\space\\
A_y = 30\sin30°,\\
B_y = 20\sin120°,\\
C_y = 10\sin225°,\\
R_y=A_y+B_y+C_y=25\text{ m}.\\\space\\
R=\sqrt{R_x^2+R_y^2}=27\text{ m},\\\space\\
\theta=\arctan\frac{R_y}{R_x}=70°."